ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 139, 2004

Bartoszynski, T., A. Nowik and T. Weiss Strongly meager sets can be quite big

237

Bell, J. P. The Hilbert series of prime PI rings  1
Fra̧czek, K. On the degree of cocycles with values in the group SU(2) 293
Franchi, C. Non-abelian sharp permutation p-groups  157
Frankl, P. and V. Rödl Strong Ramsey properties of simplices  215
Gouëzel, S. Sharp polynomial estimates for the decay of correlations  29

Green, B. Realizing deformations of curves using Lubin–Tate formal groups

139

Griffin, C. Associating curves of low genus to infinite nilpotent groups via the zeta function

67

Halbeisen, L. and E. Odell On asymptotic models in Banach spaces

253

Kaiser, C. A theorem of Nagata and Mukai/Sakai via Fourier-Mukai transform

177

Károlyi, Gy. The Erdős–Heilbronn problem in Abelian groups  349

Kohayakawa, Y., V. Rödl and P. Sissokho Embedding graphs with bounded degree in sparse pseudorandom graphs

93

Koldobsky, A., D. Ryabogin and A. Zvavitch Projections of convex bodies and the Fourier transform

361

Kozma, G. Random homeomorphisms and Fourier expansions—The pointwise behaviour

189

Kriegl, A., M. Losik and P. W. Michor Choosing roots of polynomials smoothly, II

183

Larsen, M. Word maps have large image  149

Losik, M. See A. Kriegl

Mendonça, S. and D. Zhou Maximal ergodic theorems and applications to Riemannian geometry

319

Michor, P. W. See A. Kriegl

Mudrov, A. I. Associative triples and the Yang–Baxter equation  11

Nowik, A. See T. Bartoszynski

Odell, E. See L. Halbeisen Rödl, V. See P. Frankl Rödl, V. See Y. Kohayakawa Ryabogin, D. See A. Koldobsky

Schechtman, G. Special orthogonal splittings of L_{1}^{2k}  337

Sissokho, P. See Y. Kohayakawa Weiss, T. See T. Bartoszynski Zhou, D. See S. Mendonça Zvavitch, A. See A. Koldobsky